# The Enigma of Li-Rich Giants and its Relation with Temporal Variations Observed in Radial Velocity and Stellar Activity Signals

### v1

@@ 4.1 Is the Lithium Abundance of Red Giants related to Stellar Activity?
4.2 Manifestation of Stellar Activity in RV signals and its Connection to Lithium Enhancement
@@
To further investigate these findings, we also looked at how RV variability correlates with the stellar activity indicators and whether this is linked to A(Li). All the correlations discussed here are shown in Figure 2 and listed in Appendix C. From our analysis, 22/81 stars show a strong correlation (0.4555Some of the analysed stars in our sample do not have a large number of data points. Because of this we chose to include moderate correlations which have the potential to become strong with the addition of new data. For ease of nomenclature we refer to all correlations 0.4 as strong.) between the RV variability and the BIS variability. The majority of the correlations were found to be negative. We noticed that 13/22 of the strong correlations found between RV and BIS appear for stars that are moderate and fast rotators, that is, show (e.g. de Medeiros et al. 1996; Tayar et al. 2015; Patton et al. 2023). This is likely a consequence of this diagnostic of line asymmetry losing sensitivity for slower rotators (e.g. Santos et al. 2003). There is, however, evidence in the literature (e.g. Carlberg et al. 2012; Sneden et al. 2022; Sayeed et al. 2023; Tsantaki et al. 2023) that there is a connection between Li enrichment in giants and . This is also suggested in our sample as we find that increases exponentially with A(Li) (shown in Appendix B and Table 3). Out of the 27 Li-rich stars listed in Table 2, 8 stars show strong correlations between RV and BIS.
@@
In what concerns the FWHM, 20/81 stars show strong correlations with RV measurements. The correlations are preferentially positive. 6 of the Li-rich stars in our sample show strong correlations between RV and FWHM.
@@
For the T we find that 15/81 stars show strong correlations with RV measurements. Correlations between the RV and Teff are mostly positive. Only 3 of the Li-rich stars show strong correlations between RV and T. There is a possibility that this is such because the individual spectra used to measure Teff values to look for variations may not have a signal to noise ratio high enough for this purpose.
@@
As for H16, 20/81 stars show strong correlations with RV, which are mostly negative. 9 of the Li-rich stars show strong correlations between RV and H16.
@@
Last but not least for H06, 19/81 stars show significant correlations with RV measurements, which are preferentially positive. We find that 10 of the Li-rich stars in our sample show strong correlations between RV and H06.
@@
Overall, correlations were found for stars in various evolutionary stages and of different mass ranges as well. This is evidenced in the Hertzsprung–Russell (HR) diagram Figure 3, where different symbols indicate different numbers of strong correlations found per star, and the data points are sized according to stellar mass. In general, all symbols display a range of sizes and are distributed across the HR diagram revealing there is no preferred mass or age for correlations to be found. A summary of the significant correlations discussed can be found in Tables 4 and 5. Something to note from Table 5 is that inhabitants of cluster NGC6705, in general, show higher amounts of Li than the rest of the stars in our sample. These stars are particular as they also demonstrate unusually high amounts of -elements for their young age (e.g. Casamiquela et al. 2018). Messina et al. (2010) found that the G type stars of this cluster retain higher levels of activity up to the age of the cluster. We find that, 19/28 of these stars show at least one significant correlation with one of the parameters at study and 22/28 are fast rotators (see Table 5). So in general, the stars in this cluster are magnetically active and also rotate fast which could be connected to the abnormally high levels of Li observed.
@@
In Figure 4 we show a violin plot of A(Li) as a function of the number of strong correlations observed per star. We may see that the density of points increases towards higher values of A(Li) with increasing number of correlations found. This is evidenced by the increasing width of the violins from left to right in high A(Li) regions and by the central tendency being slightly higher for stars that show 1 or more than 2 correlations between RV and stellar activity indicators. As such, strong correlations tend to appear more often for stars with higher amounts of Li in their atmospheres. Such a result also indicates that A(Li) seems to be connected with higher levels of stellar activity in our sample.
@@
To complement the correlation analysis, GLSP of RV and stellar activity indicators were also plotted for all stars. GLSP are often used when analysing whether variability observed in RV signals is caused by the existence of a companion or by stellar activity. This is because, they allow us to determine if there is periodic variability occurring at the stellar surface that is equivalent to periodic variability observed in RV. It has been shown that magnetic phenomena appearing at the stellar surface can introduce variability in RV that is at a different phase relatively to variability observed in stellar activity indicator signals (e.g. Queloz et al. 2001; Santos et al. 2014; Delgado Mena et al. 2018). As such, when looking for correlations between RV and activity proxies we may obtain a weak correlation coefficient, but this does not necessarily mean that RV variability does not originate from stellar activity. Thus, finding peaks at similar periods in RV and again in different stellar activity signals with GLSP, is also an indicator that RV variability connects with stellar activity.
@@
The results of the GLSP analysis are also presented in Tables 4 and 5. In summary, stars NGC2345No50, NGC2423No3, NGC3532No670, NGC4349No127 and NGC6705No1101 show significant peaks (above a false alarm probability (FAP) level of 1%) in their RV GLSPs which consistently appear around the same period for at least one of their stellar activity indicators. All relevant GLSPs are in Appendix D and Delgado Mena et al. (2023). All of these stars can be considered Li-rich (see Table 2) and most of them also present long-term and large amplitude RV variations which mimic the presence of planets but are of probable stellar origin (Delgado Mena et al. 2023). These stars with periodic RV variations lie on the RGB or RGB tip (they are among the most luminous in their cluster) rather than in the red clump, as found in recent works for super-Li-rich stars (Singh et al. 2019; Kumar et al. 2020). Interestingly, Jorissen et al. (2020) studying the binarity fraction of Li-rich giants found that some of the most luminous Li-rich giants show RV variability not caused by stellar companions, though no information about stellar activity is provided by the authors. In a similar way, Gonçalves et al. (2020) finds hints of RV variability in several Li-rich stars with detected longitudinal magnetic field, but the quantity of data does not allow to discern the real origin of such variability.
@@ 5 Conclusions
In summary based on the results from HARPS data, we find that the variability of BIS and FWHM and the mean and median of H06 increase exponentially with A(Li). This leads us to conclude that the stars in our sample that are more magnetically active also show higher amounts of Li in their atmosphere. We also find that stars with higher amounts of Li tend to show more correlations between RV and stellar activity indicators. The majority of stars, 19/27 (10/14 leaving aside the special cluster NGC6705), that can be considered Li-rich in this work, either show correlations of RV with at least one of the stellar activity indicators or show periodic variability in at least one of the stellar activity indicator signals with a similar period in RV. From the GLSP analysis we may also conclude that almost all the stars with strong RV variations that mimic the presence of planets are Li-rich (see also, Delgado Mena et al. 2016). This is particularly important when searching for exoplanets orbiting evolved stars, since our results show that detecting periodic RV variability in red giants with high A(Li) might be a signature of stellar activity instead of planets. It is interesting to note that Adamów et al. (2018) reported a higher than expected fraction of planetary companions around the Li-rich stars in their sample, pointing to a connection between Li enrichment and planetary presence/engulfment. However, the orbits they find are quite eccentric, something not expected if the RV were of stellar origin, and that supports the hypothesis of inner planets engulfment. The Li-rich stars with strong RV variability discussed here lie on the RGB or RGB tip, however, stars showing correlations between Li abundance and activity level are found all across the HR diagram. This might indicate that mild Li enrichment (since we do not have super-Li-rich giants in our sample) connected to stellar activity can occur at different evolutionary stages, but a large enhancement of Li is more particular and can only occur at certain evolutionary stages, such as, the He-flash (Singh et al. 2019; Kumar et al. 2020). It is still not clear then, whether the strong RV variability of mild Li-rich giants on the RGB is connected to their evolutionary stage and the possible super-Li rich phase to happen soon after (at the beginning of the He-core burning phase), or if it is a consequence of stellar activity.
@@ Acknowledgements.
I.R. and E.D.M. acknowledge the support from Fundação para a Ciência e a Tecnologia (FCT) through national funds and from FEDER through COMPETE2020 by the following grants: UIDB/04434/2020 & UIDP/04434/2020. I.R. acknowledges the support from FCT through the grant 2022.03993.PTDC. E.D.M. acknowledges the support from FCT through Stimulus FCT contract 2021.01294.CEECIND. E.D.M and J.G.S. acknowledge the support from FCT through the grant 2022.04416.PTDC.References
@@
- Bagdonas et al. (2018) Bagdonas, V., Drazdauskas, A., Tautvaišienė, G., Smiljanic, R., & Chorniy, Y. 2018, Astronomy and Astrophysics, 615, A165
@@
- Charbonnel & Balachandran (2000) Charbonnel, C. & Balachandran, S. C. 2000, arXiv preprint astro-ph/0005280
@@
- Delgado Mena et al. (2016) Delgado Mena, E., Tsantaki, M., Sousa, S., et al. 2016, Astronomy & Astrophysics, 587, A66
@@
- Frinchaboy & Majewski (2008) Frinchaboy, P. M. & Majewski, S. R. 2008, The Astronomical Journal, 136, 118
@@
- Herzog et al. (1975) Herzog, A. D., Sanders, W. L., & Seggewiss, W. 1975, Astronomy and Astrophysics Supplement Series, 19, 211
@@
- Katime Santrich et al. (2013) Katime Santrich, O. J., Pereira, C. B., & Drake, N. A. 2013, Astronomy and Astrophysics, 554, A2
- Kirby et al. (2016) Kirby, E. N., Guhathakurta, P., Zhang, A. J., et al. 2016, 819, 135
@@
- Lovis & Mayor (2007) Lovis, C. & Mayor, M. 2007, Astronomy and Astrophysics, 472, 657
- Luck (1994) Luck, R. E. 1994, The Astrophysical Journal Supplement Series, 91, 309
@@
- Mermilliod et al. (2008) Mermilliod, J. C., Mayor, M., & Udry, S. 2008, Astronomy and Astrophysics, 485, 303
@@
- Palacios et al. (2004) Palacios, A., Charbonnel, C., Talon, S., & Siess, L. 2004, arXiv e-prints, astro
- Pallavicini et al. (1992) Pallavicini, R., Randich, S., & Giampapa, M. 1992, Astronomy and Astrophysics, 253, 185
- Patton et al. (2023) Patton, R. A., Pinsonneault, M. H., Cao, L., et al. 2023, arXiv e-prints, arXiv:2303.08151
@@
- Queloz et al. (2000) Queloz, D., Mayor, M., Weber, L., et al. 2000, Astronomy and Astrophysics, 354, 99
@@
- Sayeed et al. (2023) Sayeed, M., Ness, M. K., Montet, B. T., et al. 2023, arXiv e-prints, arXiv:2306.03323
@@
- Sneden (1973) Sneden, C. A. 1973, PhD thesis, University of Texas
- Sun et al. (2022) Sun, Q., Deliyannis, C. P., Twarog, B. A., et al. 2022, Monthly Notices of the Royal Astronomical Society, 513, 5387
@@
- Udry et al. (2000) Udry, S., Mayor, M., Naef, D., et al. 2000, Astronomy and Astrophysics, 356, 590
@@
- Zboril et al. (1997) Zboril, M., Byrne, P. B., & Rolleston, W. R. J. R. 1997, Monthly Notices of the Royal Astronomical Society, 284, 685
@@
Appendix A Methods for Data Analysis
Appendix B Statistically Significant Relations
Appendix C Correlation Tables
@@
Appendix D Generalized Lomb-Scargle Periodograms

### v2

@@ 4.1 The relationship between lithium abundance and stellar activity in red giants
4.2 Manifestation of stellar activity in radial velocity signals and its connection to lithium enhancement
@@
To further investigate these findings, we also looked at how RV variability correlates with the stellar activity indicators and whether this is linked to A(Li). All the correlations discussed here are shown in Figure 2 and listed in Appendix B. From our analysis, 22/81 stars show a strong correlation (0.4888Some of the analysed stars in our sample do not have a large number of data points. Because of this, we chose to include moderate correlations that have the potential to become strong with the addition of new data. For ease of nomenclature, we refer to all correlations greater than 0.4 as strong.) between the RV variability and the BIS variability. The majority of the correlations were found to be negative. As stars age, their decreases, typically falling below 5 km/s by the end of the main-sequence (e.g. de Medeiros et al. 1996; Tayar et al. 2015; Patton et al. 2024). We noticed that 13/22 of the strong correlations found between RV and BIS appear for stars that are moderate and fast rotators (). This is likely a consequence of this diagnostic of line asymmetry losing sensitivity for slower rotators (e.g. Santos et al. 2003). There is, however, evidence in the literature (e.g. Carlberg et al. 2012; Sneden et al. 2022; Sayeed et al. 2024; Tsantaki et al. 2023) that there is a connection between Li enrichment in giants and . This is also suggested in our sample, as we find that increases exponentially with A(Li) (shown in Appendix D and Table 7). Out of the 27 Li-rich stars listed in Table 3, eight stars show strong correlations between RV and BIS.
@@
In what concerns the FWHM, 20/81 stars show strong correlations with RV measurements. The correlations are preferentially positive. Six of the Li-rich stars in our sample show strong correlations between RV and FWHM.
@@
For the T, we find that 15/81 stars show strong correlations with RV measurements. Correlations between the RV and Teff are mostly positive. Only three of the Li-rich stars show strong correlations between RV and T. There is a possibility that this the case is because the individual spectra used to measure Teff values to look for variations may not have a signal-to-noise ratio high enough for this purpose.
@@
As for H16, 20/81 stars show strong correlations with RV, and they are mostly negative. Nine of the Li-rich stars show strong correlations between RV and H16.
@@
Last but not least, for H06, 19/81 stars show significant correlations with RV measurements, and they are preferentially positive. We find that ten of the Li-rich stars in our sample show strong correlations between RV and H06.
@@
Overall, correlations were found for stars in various evolutionary stages and of different mass ranges as well. This is evidenced in the Hertzsprung–Russell (HR) diagram in Figure 3, where different symbols indicate different numbers of strong correlations found per star and the data points are sized according to stellar mass. In general, all symbols display a range of sizes and are distributed across the HR diagram, revealing there is no preferred mass or age for correlations to be found. A summary of the significant correlations discussed can be found in Tables 9 and 10. Something to note from Table 10 is that inhabitants of cluster NGC6705, in general, show higher amounts of Li than the rest of the stars in our sample. These stars are particular, as they also demonstrate unusually high amounts of -elements for their young age (e.g. Casamiquela et al. 2018). Messina et al. (2010) found that the G-type stars of this cluster retain higher levels of activity up to the age of the cluster. We find that 19/28 of these stars show at least one significant correlation with one of the parameters we studied, and 22/28 are fast rotators (see Table 10). So, in general, the stars in this cluster are magnetically active and also rotate fast, which could be connected to the abnormally high levels of Li observed.
@@
In Figure 4, we show a violin plot of A(Li) as a function of the number of strong correlations observed per star. One may see that the density of the points increases towards higher values of A(Li), with an increasing number of correlations found. This is evidenced by the increasing width of the violins from left to right in high A(Li) regions and by the central tendency being slightly higher for stars that show one or more than two correlations between RV and stellar activity indicators. As such, strong correlations tend to appear more often for stars with higher amounts of Li in their atmospheres. Such a result also indicates that A(Li) seems to be connected with higher levels of stellar activity in our sample.
@@
To complement the correlation analysis, GLSPs of RV and stellar activity indicators were also plotted for all stars. The GLSPs are often used when analysing whether variability observed in RV signals is caused by the existence of a companion or by stellar activity. This is because they make it possible to determine if there is periodic variability occurring at the stellar surface that is equivalent to the periodic variability observed in RV. It has been shown that magnetic phenomena appearing at the stellar surface can introduce variability in RV that is at a different phase relative to the variability observed in stellar activity indicator signals (e.g. Queloz et al. 2001; Santos et al. 2014; Delgado Mena et al. 2018). As such, when looking for correlations between RV and activity proxies, we may obtain a weak correlation coefficient, but this does not necessarily mean that the RV variability does not originate from stellar activity. Thus, finding peaks at similar periods in RV and again in different stellar activity signals with GLSP is also an indicator that RV variability connects with stellar activity.
@@
The results of the GLSP analysis are also presented in Tables 9 and 10. In summary, stars NGC2345No50, NGC2423No3, NGC3532No670, NGC4349No127, and NGC6705No1101 show significant peaks (above a false-alarm probability (FAP) level of 1%) in their RV GLSPs, which consistently appear around the same period for at least one of their stellar activity indicators. All relevant GLSPs are in Appendix E and Delgado Mena et al. (2023). All of these stars can be considered Li-rich (see Table 3), and most of them also present long-term and large amplitude RV variations that mimic the presence of planets but are of probable stellar origin (Delgado Mena et al. 2023). These stars with periodic RV variations lie on the RGB or RGB tip (they are among the most luminous in their cluster) rather than in the red clump, as found in recent works for super-Li-rich stars (Singh et al. 2019; Kumar et al. 2020). Interestingly, Jorissen et al. (2020), studying the binarity fraction of Li-rich giants, found that some of the most luminous Li-rich giants show RV variability not caused by stellar companions, though no information about stellar activity is provided by the authors. Similarly, Gonçalves et al. (2020) found hints of RV variability in several Li-rich stars with a detected longitudinal magnetic field, but the quantity of data does not permit one to discern the real origin of such variability.
@@ 5 Conclusions
In summary, based on the results from the HARPS data, we find that the variability of BIS and FWHM and the mean and median of H06 increase exponentially with A(Li). This leads us to conclude that the stars in our sample that are more magnetically active also show higher amounts of Li in their atmosphere. We also find that stars with higher amounts of Li tend to show more correlations between RV and stellar activity indicators. The majority of the stars, 19/27 (10/14 leaving aside the special cluster NGC6705), that can be considered Li-rich in this work either show correlations of RV with at least one of the stellar activity indicators or show periodic variability in at least one of the stellar activity indicator signals with a similar period in RV. From the GLSP analysis we may also conclude that almost all the stars with strong RV variations that mimic the presence of planets are Li-rich (see also, Delgado Mena et al. 2016). This is particularly important when searching for exoplanets orbiting evolved stars since our results show that detecting periodic RV variability in red giants with high A(Li) might be a signature of stellar activity instead of planets. It is interesting to note that Adamów et al. (2018) reported a higher than expected fraction of planetary companions around the Li-rich stars in their sample, pointing to a connection between Li enrichment and planetary presence or engulfment. However, the orbits they found are quite eccentric, something not expected if the RV is of stellar origin, and that characteristic supports the hypothesis of inner planet engulfment. The Li-rich stars with strong RV variability discussed here lie on the RGB or RGB tip. However, stars showing correlations between Li abundance and activity level are found all across the HR diagram. This might indicate that mild Li enrichment (since we do not have super Li-rich giants in our sample) connected to stellar activity can occur at different evolutionary stages, but a large enhancement of Li is more particular and can only occur at certain evolutionary stages, such as the He flash stage (Singh et al. 2019; Kumar et al. 2020). Thus, it is still not clear whether the strong RV variability of mild Li-rich giants on the RGB is connected to their evolutionary stage and the possible super Li-rich phase that happens soon after (at the beginning of the He-core burning phase) or if it is a consequence of stellar activity.
@@ Acknowledgements.
We thank the anonymous referee for the insightful comments and suggestions. I.R. and E.D.M. acknowledge the support from Fundação para a Ciência e a Tecnologia (FCT) through national funds and from FEDER through COMPETE2020 by the following grants: UIDB/04434/2020 & UIDP/04434/2020. I.R. acknowledges the support from FCT through the grant 2022.03993.PTDC. E.D.M. acknowledges the support from FCT through Stimulus FCT contract 2021.01294.CEECIND. E.D.M and J.G.S. acknowledge the support from FCT through the grant 2022.04416.PTDC. This work made use of scientific colour maps (https://www.fabiocrameri.ch/colourmaps/ Crameri et al. 2020) to mitigate visual distortion.References
@@
- Bagdonas et al. (2018) Bagdonas, V., Drazdauskas, A., Tautvaišienė, G., Smiljanic, R., & Chorniy, Y. 2018, A&A, 615, A165
@@
- Charbonnel & Balachandran (2000) Charbonnel, C. & Balachandran, S. C. 2000, A&A, 359, 563
- Crameri et al. (2020) Crameri, F., Shephard, G. E., & Heron, P. J. 2020, Nature Communications, 11, 5444
@@
- Delgado Mena et al. (2016) Delgado Mena, E., Tsantaki, M., Sousa, S. G., et al. 2016, A&A, 587, A66
@@
- Frinchaboy & Majewski (2008) Frinchaboy, P. M. & Majewski, S. R. 2008, AJ, 136, 118
@@
- Herzog et al. (1975) Herzog, A. D., Sanders, W. L., & Seggewiss, W. 1975, A&AS, 19, 211
@@
- Katime Santrich et al. (2013) Katime Santrich, O. J., Pereira, C. B., & Drake, N. A. 2013, A&A, 554, A2
- Kirby et al. (2016) Kirby, E. N., Guhathakurta, P., Zhang, A. J., et al. 2016, ApJ, 819, 135
@@
- Lovis & Mayor (2007) Lovis, C. & Mayor, M. 2007, A&A, 472, 657
- Luck (1994) Luck, R. E. 1994, ApJS, 91, 309
@@
- Mermilliod et al. (2008) Mermilliod, J. C., Mayor, M., & Udry, S. 2008, A&A, 485, 303
@@
- Palacios et al. (2006) Palacios, A., Charbonnel, C., Talon, S., & Siess, L. 2006, in Chemical Abundances and Mixing in Stars in the Milky Way and its Satellites, ed. S. Randich & L. Pasquini (Berlin, Heidelberg: Springer Berlin Heidelberg), 304–305
- Pallavicini et al. (1992) Pallavicini, R., Randich, S., & Giampapa, M. S. 1992, A&A, 253, 185
- Patton et al. (2024) Patton, R. A., Pinsonneault, M. H., Cao, L., et al. 2024, MNRAS, 528, 3232
@@
- Queloz et al. (2000) Queloz, D., Mayor, M., Weber, L., et al. 2000, A&A, 354, 99
@@
- Sayeed et al. (2024) Sayeed, M., Ness, M. K., Montet, B. T., et al. 2024, ApJ, 964, 42
@@
- Sneden (1973) Sneden, C. A. 1973, PhD thesis, The University of Texas at Austin, USA
- Sun et al. (2022) Sun, Q., Deliyannis, C. P., Twarog, B. A., et al. 2022, MNRAS, 513, 5387
@@
- Udry et al. (2000) Udry, S., Mayor, M., Naef, D., et al. 2000, A&A, 356, 590
@@
- Zboril et al. (1997) Zboril, M., Byrne, P. B., & Rolleston, W. R. J. R. 1997, MNRAS, 284, 685
@@ Appendix A Additional tables
Appendix B Correlation tables
@@
Appendix C Methods for data analysis
Appendix D Additional relations
Appendix E Generalized Lomb-Scargle periodograms